. ABC Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s expected net cash flows are as follows:


Estimated Cash Flows:
Year Project A Project B
0 -375 -575
1 -300 190
2 -200 190
3 -100 190
4 600 190
5 600 190
6 926 190
7 -200 0

If you were told that each project’s cost of capital was 12%, which project should be selected using the NPV criteria? What is each project’s IRR? What is the regular payback period for these two projects? What is the profitability index for each project if the cost of capital is 12%?
1
Expert's answer
2020-12-15T11:25:05-0500

Net present value of the projects are:

NPV = -initial cost + ACF(PVIF12%,6),

NPV(A) = $226.9.

"NPV(B) = -575 +190\u00d74.1114= 206.17."

Based on NPV creterion Project (A) should be selected because of its higher NPV.

IRR alculation of both projects using Excel function of IRR would be:

Project A = 18.64%

Project B = 23.92 %.

Regular payback periods:

Project A = 4 years + "\frac{375}{600}" = 4.625 years.

Project B ="\frac{575}{190} =" 3.03 years.

Profitability index (PI):

PI = PV cash inflow/initial costs.

"PI (A) = \frac{601.90}{375} = 1.6051."

"PI (B) = \frac{781.17}{575} = 1.3586."